OpenStudy (anonymous):

Find the exact solution (no decimal approximation) to the logarithmic equation shown: Ln(3x-5)=10

OpenStudy (anonymous):

\[3x-5=e ^{10}\] \[3x=5+e ^{10}\] \[x=5/3 + 1/3 e ^{10}\]

OpenStudy (anonymous):

Ln(3x-5)=10 10 is same as ln(e^10) so we have ln(3x-5)=ln(e^10). Since logarithm is bijective we can say 3x-5=e^10 and rest will be easy -wekku